OpenStudy (anonymous):

can you solve 3(x+14)=-4(x-7)

OpenStudy (potatoes.ramu):

3x+42= -4x+28

OpenStudy (potatoes.ramu):

7x=-14

OpenStudy (potatoes.ramu):

x=-2

OpenStudy (anonymous):

ok

OpenStudy (potatoes.ramu):

Best of luck!1 :)

OpenStudy (anonymous):

3x+42=-4x-28 +4x +4x 7x+42=-28 -42 -42 7x=-70 divide by 7 on both sides.... x=7. Theres your answer.